Find the elasticity of substitution for the following production function:Q=√LK where Q represent quantity of output produced L and K represent labour and capital used.

1
Expert's answer
2022-02-09T11:04:21-0500

"Q=\u221aLK" this is the same as "Q=L^{\frac{1}{2}}K^{\frac{1}{2}}"

"Elasticity Of Substitution=\frac{input Ratio}{MRS}"

"Inputs Ratio=\frac{K^{\frac{1}{2}}}{L^{\frac{1}{2}}}"

"MRS=\frac{-∆L}{∆K}"

"∆L=\frac{1}{2}(K^{\frac{1}{2}})"

"∆K=\frac{1}{2}(L^{\frac{1}{2}})"

"MRS=\frac{(\frac{1}{2}(K^{\frac{1}{2}})}{(\frac{1}{2}(L^\frac{1}{2})}=\frac{-K^\frac{1}{2}}{L^\frac{1}{2}}"

Therefore:

"E.S=\frac{K^{\frac{1}{2}}}{L^{\frac{1}{2}}}÷\frac{-K^{\frac{1}{2}}}{L^{\frac{1}{2}}}"

"=\frac{K^{\frac{1}{2}}}{L^{\frac{1}{2}}}×\frac{L^{\frac{1}{2}}}{-K^{\frac{1}{2}}}=-1"